OpenStudy (faman39):

A bakery charges $0.10 for a cookie that is 2 inches in diameter. If the price is proportional to the area, how much do they charge for an extra large cookie that is 10 inches in diameter? Assume that each cookie is shaped of a circle. $ 0.50 $ 1.00 $ 2.50 $ 12.50

OpenStudy (faman39):

@tkhunny can you able to help me with this ?

OpenStudy (tkhunny):

Here's the important part: "If the price is proportional to the area" Translate that Price = k*Area or just P = k*A. After that, there MUST be some given information. "$0.10 for a cookie that is 2 inches in diameter" There it is. Now what?

OpenStudy (faman39):

2.50 ?

OpenStudy (tkhunny):

Don't do that. You just went magically from a partially set up problem statement to a solution out of the air. Show your work and think it through. We have: P = k*A. and "$0.10 for a cookie that is 2 inches in diameter" Thus: 0.10 = k*pi*(2/2)^2 = k(pi) ==> k = 1/(10pi) Now what?

OpenStudy (faman39):

I m kinda confuse, you are saying multiply 2 three times then divide with 10, right?

OpenStudy (tkhunny):

No. You must read this problem very carefully. You are given the DIAMETER, but the relationship is with the AREA. You must convert. Diameter/2 = Radius. pi*Radius^2 = Area. It was mean of the question writer to use '2', since there were already two '2's in the development.

OpenStudy (faman39):

so 2 is it right?

OpenStudy (faman39):

oh ok thank you @tkhunny for helping, i will try another way

OpenStudy (tkhunny):

Once we ahve determiend 'k', we are nearly done. k = 1/(10pi) (1/10pi)*pi*(10/2)^2 = 25/10 = 2.50

OpenStudy (faman39):

It should second one?

OpenStudy (faman39):

i guess not

OpenStudy (tkhunny):

I don't understand the question. You just made me work the whole problem. It's 2.50.

OpenStudy (faman39):

I am sorry i could not get on beginning for network prob, now i went from up, where explain all, now i got. thanks so much @tkhunny !

OpenStudy (faman39):

So sorry for troubling :( i m really terrible

OpenStudy (tkhunny):

No name calling! Just spend the time it takes. You'll get it.

OpenStudy (faman39):

I am really sorry for everything :(

OpenStudy (tkhunny):

Read some material on "Direct Proportion". It's not impossible to learn it.